# Simulation of Polynomial-Time Nondeterministic Turing Machines via Universal Nondeterministic Turing Machines

### v18

@@ 5. Proving That $L_{s}\in\mathcal{NP}$
As a matter of fact, the technique to establish the desired upper bounds theorem is essentially the same as the one developed in the author’s recent work [Lin21], because the discussed issue to establish desired upper bounds is essentially the same as [Lin21]. But the work to establish such upper bounds for the first time is very difficult, and to see why, see e.g. [Lin21].
@@
On the one hand, if the input $x$ encodes a $n^{k}+k$ time-bounded nondeterministic Turing machine, then $U$ turns itself off mandatorily within
@@
In both cases we have show that for any input $w$ to $U$, there is a corresponding positive integer $i_{w}\in\mathbb{N}_{1}$ such that $U$ runs at most $|w|^{i_{w}+1}$ steps. Thus, $U$ is a nondeterministic
@@
time-bounded Turing machine for any $k\in\mathbb{N}_{1}$. By Lemma 2.7, there is a single-tape nondeterministic Turing machine $U^{\prime}$ equivalent to $U$ and $U^{\prime}$ runs within time
@@
The language $L_{s}$ is in $\mathcal{NP}$ where $L_{s}$ is accepted by the universal nondeterministic Turing machine $U$ constructed in the proof of Theorem 4.1.
@@
during the computation exceed $n^{i+1}$ steps,
which technically can be done by adding a new tape to $U$ as a counter to count up to
@@
Now, (5.1), together with (5.2) and (5.3) easily yields
@@
6. Breaking the Relativization Barrier
@@
which can be repeated arbitrarily often with different queries. If $X$ is a difficult language (say, one that cannot be decided in polynomial time, or is even undecidable), then this oracle gives the Turing machine additional power. We first quote its formal definition as follows:
@@
A nondeterministic oracle Turing machine is a nondeterministic Turing machine $M$ that has a special read-write tape we call $M$’s oracle tape and three special states $q_{query}$, $q_{yes}$, $q_{no}$. To execute $M$, we specify in addition to the input a language $X\subseteq\{0,1\}^{*}$ that is used as the oracle for $M$. Whenever during the execution $M$ enters the state $q_{query}$, the machine moves into the state $q_{yes}$ if $w\in X$ and $q_{no}$ if $w\not\in X$, where $w$ denotes the contents of the special oracle tape. Note that, regardless of the choice of $X$, a membership query to $X$ counts only as single computation step. If $M$ is an oracle machine, $X\subseteq\{0,1\}^{*}$ a language, and $x\in\{0,1\}^{*}$, then we denote the output of $M$ on input $x$ and with oracle $X$ by $M^{X}(x)$. An input $x$ is said to be accepted by a nondeterministic oracle Turing machine $M^{X}$ if there is a computation path of $M^{X}$ on input $x$ leading to the accepting state of $M^{X}$.
@@
The set of languages decided by nondeterministic oracle Turing machines with oracle $X$ within time $t(n)$ by the condition of “there exists at least one accepting path” is denoted by ${\rm NTIME}^{X}[t(n)]$. Thus,
@@
Let $M^{A}$ be a polynomial-time nondeterministic oracle Turing machine with oracle $A$. Since $M^{A}$ can be regarded as both a polynomial-time nondeterministic oracle Turing machine working under the condition of “there exists at least one accepting path” and a ${\rm co}\mathcal{NP}^{A}$ machine working under the condition of “there exists no accepting path”, we also strictly distinguish, in the following context, between these two situations in which the use of symbols. Namely, we use the notation
@@
In a similar way to define the complexity class ${\rm coNTIME}[t(n)]$ in Section 2, we can define the complexity class ${\rm coNTIME}^{X}[t(n)]$, based on the “there exists no accepting path” condition, as follows.
@@
The set of languages decided by nondeterministic oracle Turing machines with oracle $X$ within time $t(n)$ by the condition of “there exists no accepting path” is denoted by ${\rm coNTIME}^{X}[t(n)]$, i.e., a $t(n)$ time-bounded nondeterministic oracle Turing machine with oracle $X$ accepts a string $w$ if and only if there exists no accepting path on input $w$. Thus,
@@
In the following context, we will assume the notion of ${\rm co}\mathcal{P}^{X}$ machines, which can be defined in a similar way to that of Definition 6.4, and is omitted here.
@@
Note that, the result of Baker, Gill, and Solovay [BGS75] also implies that for the same oracle $O$,
@@
because if $\mathcal{P}^{O}=\mathcal{NP}^{O}$, then ${\rm co}\mathcal{P}^{O}={\rm co}\mathcal{NP}^{O}$. Further by
@@
If $\mathcal{P}^{O}=\mathcal{NP}^{O}={\rm co}\mathcal{NP}^{O}$, and assume that the set of all ${\rm co}\mathcal{P}^{O}$ machines is enumerable, then simulation techniques are unable to create a language $L_{s}^{O}\notin{\rm co}\mathcal{NP}^{O}$, thus unable to separate $\mathcal{NP}^{O}$ from ${\rm co}\mathcal{NP}^{O}$.
@@
First, note that $\mathcal{P}^{O}=\mathcal{NP}^{O}$, thus we have ${\rm co}\mathcal{P}^{O}={\rm co}\mathcal{NP}^{O}$. Clearly, in this scenario, to separate $\mathcal{NP}^{O}$ from ${\rm co}\mathcal{NP}^{O}$ is equivalent to separating $\mathcal{P}^{O}$ from ${\rm co}\mathcal{P}^{O}$. Moreover, if the set of all ${\rm co}\mathcal{P}^{O}$ machines is not enumerable, then of course we can not apply simulation techniques. Thus, we assume that the set of ${\rm co}\mathcal{P}^{O}$ machines is enumerable, as the premise of the theorem states. Since we have assumed that the set of all ${\rm co}\mathcal{P}^{O}$ machines is enumerable, the rest of the proof is similar to the Remark 4.2. Namely, the simulation technique is unable to create a language not in ${\rm co}\mathcal{P}^{O}$ via a universal deterministic oracle Turing machine $U^{O}$, because a polynomial-time deterministic oracle Turing machine that works under the “there exists no accepting path” condition is still a polynomial-time deterministic oracle Turing machine. In short, the “there exists at least one accepting path” condition and the “there exists no accepting path” condition for deterministic oracle Turing machines are symmetrical. Specifically, for an arbitrary deterministic oracle Turing machine $M^{O}$, if $M^{O}$ accepts the input $w$, then there is only one accepting path, and if $M^{O}$ rejects the input $w$, then there is also only one rejecting path. This is different from nondeterministic oracle Turing machines. If we apply the same techniques, i.e., the simulation techniques, to create a language not in ${\rm co}\mathcal{P}^{O}$, then the universal deterministic oracle Turing machine $U^{O}$ (it runs within time $O(n^{k})$ for any $k\in\mathbb{N}_{1}$) needed to be constructed accepts the language $L(U^{O})\in\mathcal{P}^{O}$ (the proof is similar to that of Theorem 6.4 below). But, by the symmetrical properties between the “there exists at least one accepting path” condition and the “there exists no accepting path” condition with respect to deterministic oracle Turing machines, we clearly have that $L(U^{O})\in{\rm co}\mathcal{P}^{O}$. Thus, we are unable to claim that $L(U^{O})\notin{\rm co}\mathcal{P}^{O}$.
@@
If $\mathcal{P}^{A}\neq\mathcal{NP}^{A}$ and suppose that the set $\mathcal{T}^{A}$ of all ${\rm co}\mathcal{NP}^{A}$ machines with oracle $A$ is enumerable and the enumeration is $e^{\prime}$, and we have the above rational assumptions. Then, there exists a language $L_{s}^{A}$ accepted by a universal nondeterministic oracle Turing machine $U^{A}$ with oracle $A$, but not by any ${\rm co}\mathcal{NP}^{A}$ machines with oracle $A$. Namely, there is a language $L_{s}^{A}$ such that
@@
in a similar way to that of Theorem 4.1. For convenience of the reader and for clarity, we proceed as follows: let $U^{A}$ be a five-tape universal nondeterministic oracle Turing machine that operates as follows on an input string $x$ of length of $n$:
@@
    $U^{A}$ decodes the tuple encoded by $x$. If $x$ is not the encoding of some ${\rm co}\mathcal{NP}^{A}$ machine $M^{A}_{j}$ with oracle $A$ for some $j$ (by Definition 6.4, equivalently, $x$ can not be decoded into some polynomial-time nondeterministic oracle Turing machine $M^{A}_{j}$ for some $j$, i.e., $x$ represents the trivial oracle Turing machine with an empty next-move function) then GOTO $6$, else determines $t$, the number of tape symbols used by $M^{A}_{j}$; $s$, its number of states; and $k$, its order. [^4] We suppose that the order $110^{k}$ 1 of $M^{A}_{j}$ for some $k\in\mathbb{N}_{1}$ is also encoded into the binary string representing ${\rm co}\mathcal{NP}^{A}$ machine $M^{A}_{j}$, which is similar to Section 3. The third tape of $U^{A}$ can be used as “scratch” memory to calculate $t$.
@@
    cells each, the blocks being separated by single cell holding a marker ##\##, i.e., there are
@@
    On tape $4$, $U^{A}$ reads and writes the contents of oracle tape of $M^{A}_{j}$. That is, tape $4$ is the oracle tape of $U^{A}$ which is used to simulate the oracle tape of $M^{A}_{j}$.
@@
    time and output its answer, i.e., accepting if $M^{A}_{j}$ accepts and rejecting if $M^{A}_{j}$ rejects.”
@@
    Since $x$ is encoding of the trivial oracle Turing machine with an empty next-move function. Then $U^{A}$ sets up a block of
@@
We claim that $L^{A}_{s}\notin{\rm co}\mathcal{NP}^{A}$. Indeed, suppose for the sake of contradiction that $L^{A}_{s}$ is decided by some $n^{k}+k$ time-bounded ${\rm co}\mathcal{NP}^{A}$ machine $M^{A}_{i}$. Then, we may assume that $M^{A}_{i}$ is a two-tape (with an input tape and an oracle tape) $T(n)=n^{k}+k$ time-bounded ${\rm co}\mathcal{NP}^{A}$ machine. By Definition 6.4, we can regard it as an $n^{k}+k$ time-bounded nondeterministic oracle Turing machine with oracle $A$, and we can further let $U^{A}$ on input $1^{n-m}\langle M^{A}_{i}\rangle$ simulate $M^{A}_{i}$ on input $1^{n-m}\langle M^{A}_{i}\rangle$. Let $M^{A}_{i}$ have $s$ states and $t$ tape symbols and the shortest length of $M^{A}_{i}$ is $m$, i.e,
@@
But by the relation between the languages $L(M^{A}_{i})$ accepted by polynomial-time nondeterministic orale Turing machine $M^{A}_{i}$ with oracle $A$ and $\overline{L}(M^{A}_{i})$ accepted by ${\rm co}\mathcal{NP}^{A}$ machine $M^{A}_{i}$, we have
@@
$1^{|w|-m}\langle M^{A}_{i}\rangle\notin L^{A}_{s}$
@@
The proof of this theorem is similar to that of Theorem 5.2. Namely, the first step is to define the family of languages
@@
Now, we are naturally at a right point to present the proof of Theorem 1.2 as follows: Proof of Theorem 1.2. Note that the statement of Theorem 1.2 is precisely the same as Theorem 6.6, thereby we have done the proof.

### v19

@@ 5. Proving That $L_{s}\in\mathcal{NP}$
As a matter of fact, the technique to establish the desired upper bounds theorem is essentially the same as the one developed in the author’s recent work [Lin21], because the discussed issue to establish desired upper bounds is essentially the same as [Lin21]. But the work to establish such upper bounds for the first time is very difficult, and to see why, see e.g., [Lin21].
@@
On the one hand, if the input $x$ encodes an $n^{k}+k$ time-bounded nondeterministic Turing machine, then $U$ turns itself off mandatorily within
@@
In both cases we have shown that for any input $w$ to $U$, there is a corresponding positive integer $i_{w}\in\mathbb{N}_{1}$ such that $U$ runs at most $|w|^{i_{w}+1}$ steps. Thus, $U$ is a nondeterministic
@@
time-bounded Turing machine for any $k\in\mathbb{N}_{1}$. By Lemma 2.7, there is a single-tape nondeterministic Turing machine $U^{\prime}$ equivalent to $U$, and $U^{\prime}$ runs within time
@@
The language $L_{s}$ is in $\mathcal{NP}$, where $L_{s}$ is accepted by the universal nondeterministic Turing machine $U$ constructed in the proof of Theorem 4.1.
@@
during the computation exceed $n^{i+1}$ steps.
Note that the above definition of language $L_{s}^{i}$ technically can be done by adding a new tape to $U$ as a counter to count up to
@@
Now, (5.1), together with (5.2) and (5.3), easily yields
@@
Remark 5.5.
From Theorem 5.1 we know that for any input $w\in\{0,1\}^{*}$, there exists a positive integer $i_{w}$ such that $U$ runs at most $|w|^{i_{w}+1}$ steps. This, together with (5.4), more clearly implies that
$L_{s}=\bigcup_{n\in\mathbb{N}_{1}}L_{s}^{n}.$
Namely, (5.1) can be deduced from the conclusion of Theorem 5.1 and (5.4).
6. Breaking the “Relativization Barrier”
@@
which can be repeated arbitrarily often with different queries. If $X$ is a difficult language (say, one that cannot be decided in polynomial time or is even undecidable), then this oracle gives the Turing machine additional power. We first quote its formal definition as follows:
@@
A nondeterministic oracle Turing machine is a nondeterministic Turing machine $M$ that has a special read-write tape we call $M$’s oracle tape and three special states $q_{query}$, $q_{yes}$, and $q_{no}$. To execute $M$, we specify in addition to the input a language $X\subseteq\{0,1\}^{*}$ that is used as the oracle for $M$. Whenever during the execution $M$ enters the state $q_{query}$, the machine moves into the state $q_{yes}$ if $w\in X$ and $q_{no}$ if $w\not\in X$, where $w$ denotes the contents of the special oracle tape. Note that, regardless of the choice of $X$, a membership query to $X$ counts only as a single computation step. If $M$ is an oracle machine, $X\subseteq\{0,1\}^{*}$ a language, and $x\in\{0,1\}^{*}$, then we denote the output of $M$ on input $x$ and with oracle $X$ by $M^{X}(x)$. An input $x$ is said to be accepted by a nondeterministic oracle Turing machine $M^{X}$ if there is a computation path of $M^{X}$ on input $x$ leading to the accepting state of $M^{X}$.
@@
The set of languages decided by all nondeterministic oracle Turing machines with oracle $X$ within time $t(n)$ by the condition of “there exists at least one accepting path” is denoted by ${\rm NTIME}^{X}[t(n)]$. Thus,
@@
Let $M^{A}$ be a polynomial-time nondeterministic oracle Turing machine with oracle $A$. Since $M^{A}$ can be regarded as both a polynomial-time nondeterministic oracle Turing machine working under the condition of “there exists at least one accepting path” and a ${\rm co}\mathcal{NP}^{A}$ machine working under the condition of “there exists no accepting path,” we also strictly distinguish, in the following context, between these two situations in which the use of symbols. Namely, we use the notation
@@
In a similar way to defining the complexity class ${\rm coNTIME}[t(n)]$ in Section 2, we can define the complexity class ${\rm coNTIME}^{X}[t(n)]$, based on the “there exists no accepting path” condition, as follows.
@@
The set of languages decided by all nondeterministic oracle Turing machines with oracle $X$ within time $t(n)$ by the condition of “there exists no accepting path” is denoted by ${\rm coNTIME}^{X}[t(n)]$, i.e., a $t(n)$ time-bounded nondeterministic oracle Turing machine with oracle $X$ accepts a string $w$ if and only if there exists no accepting path on input $w$. Thus,
@@
In the following context, we will assume the notion of ${\rm co}\mathcal{P}^{X}$ machines, which can be defined in a similar way to that of Definition 6.4 and is omitted here.
@@
Note that the result of Baker, Gill, and Solovay [BGS75] also implies that for the same oracle $O$,
@@
because if $\mathcal{P}^{O}=\mathcal{NP}^{O}$, then ${\rm co}\mathcal{P}^{O}={\rm co}\mathcal{NP}^{O}$. Further, by
@@
If $\mathcal{P}^{O}=\mathcal{NP}^{O}={\rm co}\mathcal{NP}^{O}$, and we assume that the set of all ${\rm co}\mathcal{P}^{O}$ machines is enumerable, then simulation techniques are unable to create a language $L_{s}^{O}\notin{\rm co}\mathcal{NP}^{O}$, thus unable to separate $\mathcal{NP}^{O}$ from ${\rm co}\mathcal{NP}^{O}$.
@@
First, note that $\mathcal{P}^{O}=\mathcal{NP}^{O}$, thus we have ${\rm co}\mathcal{P}^{O}={\rm co}\mathcal{NP}^{O}$. Clearly, in this scenario, to separate $\mathcal{NP}^{O}$ from ${\rm co}\mathcal{NP}^{O}$ is equivalent to separating $\mathcal{P}^{O}$ from ${\rm co}\mathcal{P}^{O}$. Moreover, if the set of all ${\rm co}\mathcal{P}^{O}$ machines is not enumerable, then of course we can not apply simulation techniques. Thus, we assume that the set of ${\rm co}\mathcal{P}^{O}$ machines is enumerable, as the premise of the theorem states. Since we have assumed that the set of all ${\rm co}\mathcal{P}^{O}$ machines is enumerable, the rest of the proof is similar to Remark 4.2. Namely, the simulation technique is unable to create a language not in ${\rm co}\mathcal{P}^{O}$ via a universal deterministic oracle Turing machine $U^{O}$, because a polynomial-time deterministic oracle Turing machine that works under the “there exists no accepting path” condition is still a polynomial-time deterministic oracle Turing machine. In short, the “there exists at least one accepting path” condition and the “there exists no accepting path” condition for deterministic oracle Turing machines are symmetrical. Specifically, for an arbitrary deterministic oracle Turing machine $M^{O}$, if $M^{O}$ accepts the input $w$, then there is only one accepting path, and if $M^{O}$ rejects the input $w$, then there is also only one rejecting path. This is different from nondeterministic oracle Turing machines. If we apply the same techniques, i.e., the simulation techniques, to create a language not in ${\rm co}\mathcal{P}^{O}$, then the universal deterministic oracle Turing machine $U^{O}$ (it runs within time $O(n^{k})$ for any $k\in\mathbb{N}_{1}$) needed to be constructed accepts the language $L(U^{O})\in\mathcal{P}^{O}$ (the proof is similar to that of Theorem 6.4 below). But, by the symmetrical properties between the “there exists at least one accepting path” condition and the “there exists no accepting path” condition with respect to deterministic oracle Turing machines, we clearly have that $L(U^{O})\in{\rm co}\mathcal{P}^{O}$. Thus, we are unable to claim that $L(U^{O})\notin{\rm co}\mathcal{P}^{O}$.
@@
If $\mathcal{P}^{A}\neq\mathcal{NP}^{A}$ and we suppose that the set $\mathcal{T}^{A}$ of all ${\rm co}\mathcal{NP}^{A}$ machines with oracle $A$ is enumerable and the enumeration is $e^{\prime}$, and we have the above rational assumptions. Then, there exists a language $L_{s}^{A}$ accepted by a universal nondeterministic oracle Turing machine $U^{A}$ with oracle $A$, but not by any ${\rm co}\mathcal{NP}^{A}$ machines with oracle $A$. Namely, there is a language $L_{s}^{A}$ such that
@@
in a similar way to that of Theorem 4.1. For the convenience of the reader and for clarity, we proceed as follows: Let $U^{A}$ be a five-tape universal nondeterministic oracle Turing machine that operates as follows on an input string $x$ of length $n$:
@@
    $U^{A}$ decodes the tuple encoded by $x$. If $x$ is not the encoding of some ${\rm co}\mathcal{NP}^{A}$ machine $M^{A}_{j}$ with oracle $A$ for some $j$ (by Definition 6.4, equivalently, $x$ can not be decoded into some polynomial-time nondeterministic oracle Turing machine $M^{A}_{j}$ for some $j$, i.e., $x$ represents the trivial oracle Turing machine with an empty next-move function), then GOTO $6$, else determine $t$, the number of tape symbols used by $M^{A}_{j}$; $s$, its number of states; and $k$, its order.  [^5] We suppose that the order $110^{k}$ 1 of $M^{A}_{j}$ for some $k\in\mathbb{N}_{1}$ is also encoded into the binary string representing ${\rm co}\mathcal{NP}^{A}$ machine $M^{A}_{j}$, which is similar to Section 3. The third tape of $U^{A}$ can be used as “scratch” memory to calculate $t$.
@@
    cells each, the blocks being separated by a single cell holding a marker ##\##, i.e., there are
@@
    On tape $4$, $U^{A}$ reads and writes the contents of the oracle tape of $M^{A}_{j}$. That is, tape $4$ is the oracle tape of $U^{A}$, which is used to simulate the oracle tape of $M^{A}_{j}$.
@@
    time and outputs its answer, i.e., accepting if $M^{A}_{j}$ accepts and rejecting if $M^{A}_{j}$ rejects.”
@@
    Since $x$ is an encoding of the trivial oracle Turing machine with an empty next-move function. Then $U^{A}$ sets up a block of
@@
We claim that $L^{A}_{s}\notin{\rm co}\mathcal{NP}^{A}$. Indeed, suppose for the sake of contradiction that $L^{A}_{s}$ is decided by some $n^{k}+k$ time-bounded ${\rm co}\mathcal{NP}^{A}$ machine $M^{A}_{i}$. Then, we may assume that $M^{A}_{i}$ is a two-tape (with an input tape and an oracle tape) $T(n)=n^{k}+k$ time-bounded ${\rm co}\mathcal{NP}^{A}$ machine. By Definition 6.4, we can regard it as an $n^{k}+k$ time-bounded nondeterministic oracle Turing machine with oracle $A$, and we can further let $U^{A}$ on input $1^{n-m}\langle M^{A}_{i}\rangle$ simulate $M^{A}_{i}$ on input $1^{n-m}\langle M^{A}_{i}\rangle$. Let $M^{A}_{i}$ have $s$ states and $t$ tape symbols, and the shortest length of $M^{A}_{i}$ is $m$, i.e.,
@@
But by the relation between the languages $L(M^{A}_{i})$ accepted by polynomial-time nondeterministic oracle Turing machine $M^{A}_{i}$ with oracle $A$ and $\overline{L}(M^{A}_{i})$ accepted by ${\rm co}\mathcal{NP}^{A}$ machine $M^{A}_{i}$, we have
@@
$1^{|w|-m}\langle M^{A}_{i}\rangle\notin L^{A}_{s},$
@@
The proof of this theorem is similar to that of Theorem 5.2. Namely, the first step is to define the family of languages:
@@
Now, we are naturally at a right point to present the proof of Theorem 1.2 as follows: Proof of Theorem 1.2. Note that the statement of Theorem 1.2 is precisely the same as Theorem 6.6; therefore, we have done the proof.